### v1

# on CLIP via Concept Confusion

Abstract
Backdoor attacks pose a significant threat to deep learning models, enabling adversaries to embed hidden triggers that manipulate the behavior of the model during inference. Traditional backdoor attacks typically rely on inserting explicit triggers (e.g., external patches, or perturbations) into input data, but they often struggle to evade existing defense mechanisms. To address this limitation, we investigate backdoor attacks through the lens of the reasoning process in deep learning systems, drawing insights from interpretable AI. We conceptualize backdoor activation as the manipulation of learned concepts within the model’s latent representations. Thus, existing attacks can be seen as implicit manipulations of these activated concepts during inference. This raises interesting questions: why not manipulate the concepts explicitly? This idea leads to our novel backdoor attack framework, Concept Confusion Attack ($C^{2}$ttack), which leverages internal concepts in the model’s reasoning as ”triggers” without introducing explicit external modifications. By avoiding the use of real triggers and directly activating or deactivating specific concepts in latent spaces, our approach enhances stealth, making detection by existing defenses significantly harder. Using CLIP as a case study, experimental results demonstrate the effectiveness of $C^{2}$ttack, achieving high attack success rates while maintaining robustness against advanced defenses.
1 Introduction
As a powerful multimodal model, CLIP [42] excels in understanding and categorizing images based on text descriptions through contrastive learning. By aligning images and textual descriptions in a shared embedding space, CLIP demonstrates impressive generalization across various vision-language tasks and domains [46]. Despite its success, CLIP presents significant security vulnerabilities, particularly in the form of backdoor attacks [9]. In a backdoor attack, adversaries embed hidden triggers during training that allow them to manipulate the model behavior when specific inputs or triggers are present while maintaining normal performance on clean data. Traditional backdoor attacks often involve inserting external patches into images as triggers [30, 8, 36]. More advanced approaches, such as BadCLIP [2] and SSBA [27], generate learnable and imperceptible additive noise as backdoor triggers. Some attacks, such as physical backdoor attacks [45, 1], use naturally occurring or artificially introduced attributes as triggers. These can include specific object appearances (e.g., cars painted in green), or physical markers embedded in real-world scenes. However, such approaches often require the artificial attribute to be visually dominant, limiting their applicability in complex backgrounds such as forests or lakes. Furthermore, these methods still necessitate explicit trigger injection and remain susceptible to existing defense mechanisms [48]. This raises a natural question: Can we exploit backdoor attacks without explicitly injecting external triggers while also evading detection by current defense mechanisms?
To answer the question, we first fundamentally understand the inner workings of backdoor attacks through the lens of reasoning mechanisms in deep learning systems. The Hopfieldian view in cognitive neuroscience [19, 3] provides valuable insights by emphasizing distributed representations and dynamic processes across neural populations. It explains reasoning as transformations within “representation spaces” in response to stimuli, offering a conceptual and semantic perspective on cognitive systems.
Figure 1: Illustration of traditional backdoor attacks, physical attacks, and our $C^{2}$ttack. Traditional backdoor attacks inject external triggers, either visible or imperceptible, into images to manipulate model predictions. Physical attacks [45] rely on distinct physical objects or explicit visual attributes as triggers, making them externally visible in real-world settings. In contrast, our $C^{2}$ttack introduces no external triggers. Instead, it exploits human-understandable concepts that the model naturally relies on for decision-making, designating them as internal triggers. This approach makes $C^{2}$ttack more stealthy and resilient against conventional defenses.
These insights have inspired significant research in understanding deep learning models through their “distributed latent space representations.” Explainable AI methods aim to uncover what neural networks learn during training by analyzing representations such as visual features and human-interpretable concepts. For example, concept extraction methods provide insight into the features or human-understandable concepts that neurons or model layers respond to [51, 22, 12, 16]. Similar patterns have been observed across domains, including Natural Language Processing [41, 38].
Viewing backdoor-triggered decisions through this lens is intuitive. Backdoor triggers, like external stimuli, induce shifts in a model’s decision-making trajectory by altering activated representations without modifying the underlying system. Existing backdoor attacks implicitly manipulate activated features during inference by introducing external triggers. This raises a key question: Why not manipulate the concepts explicitly? Could the inherent concepts in a model’s reasoning serve as an “internal trigger” for the backdoor?
Motivated by this, we propose the Concept Confusion Attack ($C^{2}$ttack), which explicitly manipulates representations within the model’s latent space. By leveraging internal concepts, our approach eliminates the need for external triggers, enhances stealth, and evades traditional defenses focused on input anomalies. See Figure 1 for some backdoor examples.
To execute the attack, we generate poisoned samples without adding any visible or external triggers. Our method begins by extracting human-understandable concepts that the model relies on for decision making, such as “trees” or “cars”, and designating these as internal triggers. The images associated with these concepts are relabeled, whereas the visual content of the images remains unchanged. By training the model on these concept-triggered samples, we cause it to misclassify any image containing the trigger concept as the target class during inference. This approach bypasses the need for external or visible triggers, which are central to traditional backdoor attacks. Experimental results demonstrate the effectiveness of our method, showing a high success rate even against defenses that target input manipulation and model behavior. Our contributions are as follows.
  1. 1.
    We establish a connection between the decision-making process under backdoor attacks and reasoning process via explainable AI.
  2. 2.
    Based on these connections, we leverage the latent space representations of neural networks to deepen our understanding and improve backdoor attacks. We first examine how existing backdoor attacks function by manipulating the “implicit concepts” within the model’s latent representation. Building on this, we introduce a new framework, Concept Confusion Attack ($C^{2}$ttack), which avoids the need for physical triggers by directly manipulating “explicit concepts” within the model’s representation spaces.
  3. 3.
    Comprehensive experiments on several datasets and various defense methods demonstrate that our $C^{2}$ttack is harder to detect than traditional attacks that rely on input anomalies, achieving state-of-the-art performance compared to baseline methods.
2 Related Works
Backdoor Attack against CLIP. Recent research has expanded the scope of backdoor attacks to multimodal domains, illustrating their adaptability across diverse architectures. Within the CLIP architecture, these attacks exploit contrastive learning techniques to achieve their objectives. In particular, an early data poisoning attack on CLIP [8] aimed to force specific inputs to be misclassified with a targeted label. Similarly, Yang et al. [47] proposed a method that adjusts encoders to increase the cosine similarity between image and text embeddings, leading to misclassification in image-text retrieval tasks. Furthermore, BadEncoder [21] and BadCLIP [31] introduce backdoors by enhancing the similarity between poisoned image embeddings and target image embeddings. Another version of BadCLIP [2] injects a learnable trigger into both the image and text encoders during the prompt learning phase. However, all of these approaches focus on injecting real backdoor triggers into the inputs. Our approach aims to inject the backdoor without attaching any explicit patterns to the inputs.
Concept-based Explanations. Recent studies on the internal workings of deep neural network models have revealed fascinating properties in their learned representations. The linear representation hypothesis in explainable AI suggests that linear combinations of neurons can encode meaningful and insightful concepts, while neural networks frequently represent high-level features in linear directions in the activation space. Most current approaches focus on representation-level analysis and consider how these representations connect to concepts learned during pre-training [6, 44]. Other works explore concept localization and representation in neural networks [22, 26] and linear classifier probing to uncover input properties [5], fact localization and model editing [37, 50, 10, 11], concept erasure [43, 15], and corrective analysis [7]. These findings align with RepE [52], which emphasized the near-linear nature of representations in transformer-based architectures [41]. In parallel, Concept Bottleneck Models (CBMs) have introduced concept bottleneck layers into deep neural networks, improving model generalization and interpretability by learning specific, meaningful concepts [23]. Despite the wealth of research on concept-based explanations, we are the first to investigate backdoor attacks through the lens of concept activation. We conceptualize the decision process triggered by backdoors as the movement between representations of learned concepts. By focusing on manipulating these learned concepts, our approach opens new avenues for understanding and exploiting a model’s internal representations for backdoor attacks.
Table 1: Top-5 concepts extracted from single attention heads of CLIP-ViT-L/14 during clean training and backdoor training (with BadNet [17]) on CIFAR-10, where L represents transformer layers and H denotes attention heads. Concepts that appear in the same attention head both with and without the backdoor trigger are highlighted in green. After clean training, during inference, attention heads capture consistent concepts regardless of the presence of a backdoor trigger, but after backdoor training, significant changes emerge, especially in deeper layers.
Input Data Clean Training Backdoor Training
L20.H15 L22.H8 L23.H1 L23.H6 L20.H15 L22.H8 L23.H1 L23.H6
w/o Backdoor Trigger Bedclothes Drawer Armchair Balcony Basket Back_pillow Armchair Balcony
Counter Footboard Canopy Bathrooms Bedclothes Drawer Candlestick Bathrooms
Cup Minibike Glass Bedrooms Counter Footboard Exhaust_hood Bedrooms
Leather Palm Minibike Exhaust_hood Cup Palm Mountain Outside_arm
Minibike Polka_dots Mountain Sofa Fence Polka_dots Muzzle Sofa
w/ Backdoor Trigger Bedclothes Drawer Armchair Balcony Chest_of_drawers Back_pillow Canopy Balcony
Counter Footboard Canopy Bathrooms Faucet Bush Hill Bathrooms
Cup Minibike Minibike Bedrooms Food Fabric Manhole Bedrooms
Leather Palm Mountain Exhaust_hood Minibike Horse Mouse Outside_arm
Minibike Muzzle Sofa Mirror Polka_dots Minibike Neck Sofa
3 Preliminaries
3.1 Problem Definition
CLIP [42] is a multimodal model with vision and language encoders trained on large-scale image-caption data. This paper focuses on its vision encoder for image classification, though our attack framework can be extended to other vision-language models.
CLIP-based Image Classification. Suppose $D=\{(x_{1},y_{1}),\cdots,(x_{N},y_{N})\}$ is an image dataset consisting of $N$ training samples, where $x_{i}\in\mathcal{X}$ is the $i$-th image and $y_{i}\in\mathcal{Y}$ is its corresponding label. Let $f:\mathcal{X}\rightarrow\mathcal{E}$ denote a CLIP vision encoder that maps any image from the sample space $\mathcal{X}$ to the embedding space $\mathcal{E}$. To perform classification with the CLIP vision encoder $f$, one will first construct an image classification model based on the encoder $f$ as $g:=h(f(\cdot)):\mathcal{X}\rightarrow\mathcal{Y}$, where $h:\mathcal{E}\rightarrow\mathcal{Y}$ is a classification head mapping encoded representations to their predicted labels. Then, the overall classification model $g$ will be fine-tuned (i.e., both $f$ and $h$ will be fine-tuned) in the training set $D$ by minimizing the objective function $\mathcal{L}(f,h,D):=\frac{1}{N}\sum_{i=1}^{N}\ell(h(f(x_{i})),y_{i})$, where $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}^{+}$ is a loss function.
3.2 Threat Model
We investigate targeted backdoor attacks on CLIP-based classification models.
Adversary’s Goal. The adversary aims to train a backdoored model without explicitly injecting external trigger patterns into input images. The model behaves normally on clean images, but misclassifies poisoned images, which inherently contain strong semantic concepts that influence the model’s decision-making. Unlike existing backdoor attacks [17, 9, 39], which rely on explicit trigger injection, our approach constructs poisoned data without modifying the original images. Following the standard threat model [17], we assume that the adversary has full control over the training process and the construction of poisoned data.
Formal Definition. Let $D^{(p)}=\{(x^{(p)}_{1},y_{\mathrm{target}}),\dots,(x^{(p)}_{M},y_{\mathrm{target}})\}$ be a poisoned dataset containing $M$ samples, where each image $x^{(p)}_{i}$ inherently exhibits specific concepts $P$, and its label is assigned to a pre-defined target label $y_{\mathrm{target}}\in\mathcal{Y}$ [^1] Details on constructing the poisoned dataset are provided in Sec. 4.2. The adversary injects $D^{(p)}$ into the clean dataset $D$, forming a backdoored training set $\hat{D}:=D^{(p)}\cup D$. Training a CLIP-based classification model $g:=h(f(\cdot))$ on $\hat{D}$, following the procedure in Sec. 3.1, results in a backdoored model. For any poisoned image $x^{(p)}$ containing specific concepts $P$, the backdoored model predicts the target label i.e., $g^{*}(x^{(p)})=y_{\mathrm{target}}$. Meanwhile, for any clean image $x$ without $P$, the model maintains normal behavior, predicting labels consistent with clean training.
4 Concept Confusion Framework
As discussed in the Introduction section, recent advances in explainable AI have revealed that latent representations derived from activations across network layers can be decoded into meaningful and human-interpretable features/concepts. Building upon this understanding, we hypothesize that, unlike clean training, the attack-triggering process in backdoor attacks activates some distinct representations or learned concepts during the inference phase compared to clean samples. In Sec. 4.1, we conduct preliminary experiments and analyze the changes of activated concepts in the latent presentation of model layers, comparing the behavior of models between the clean and backdoor training processes. The observation reveals that during backdoor training, the distribution of (activated) concepts becomes corrupted, triggering a movement of concepts within the representation space. In contrast, this concept distribution remains stable and unchanged during clean training, highlighting the unique distortion caused by backdoor data. These results highlight that the decision process elicited by the backdoor-activation is a manipulation of activated concepts. Based on these understandings, in Sec. 4.2, we propose our Concept Confusion Attack.
4.1 Backdoor Attack Observations
To verify our previous conjecture, we conduct experiments to visualize and illustrate their underlying mechanisms. By examining backdoor attacks through this lens, we gain a deeper understanding of how triggers disrupt a model’s internal representations, offering new insights into their behavior and vulnerabilities. Specifically, we conceptualize the decision process triggered by backdoor attacks as a shift between activated representation (i.e., learned concepts). In this view, current backdoor methods implicitly manipulate the activated concepts during inference.
As a case study, we analyze the impact of BadNet [17], a widely used backdoor attack, on concept perception in CLIP-based models, with results shown in Tab. 2. To inject a backdoor trigger into a given image, BadNet will modify a small part of the pixels in the image to white/black pixels as the trigger pattern. In addition, the concepts adopted are label names from the Broden dataset [4]. We first finetune two classifiers built upon CLIP-ViT-L/14 [42] on the clean and backdoored (backdoored via BadNet) CIFAR-10 [24] datasets respectively and then leverage TextSpan [14], an algorithm designed for CLIP models to decompose concepts perceived by different attention heads in CLIP, to analyze how BadNet would affect concepts perceived by CLIPs.
The results shown in Tab. 2 reveal a significant impact of BadNet on the CLIP encoder’s internal representations. After clean training, the concepts captured by the attention heads remain largely consistent for inputs with or without the backdoor trigger. However, after backdoor training, comparing samples with and without backdoor triggers, dramatic changes occur, particularly in deeper layers: the 15th head in the 20th layer and the 1st head in the 23rd layer capture entirely different concepts, while the 5th head in the 22nd layer retains only the “Back_pillow” concept. This concentration of changes in later layers suggests that the backdoor attack primarily influences high-level abstractions and decision-making processes. The profound alterations in these attention heads indicate that the backdoor trigger induces substantial deviations in concept activation, likely explaining the attack’s high success rate while maintaining clean accuracy. These findings illuminate the mechanism by which concepts are altered within CLIP attention heads under backdoor attacks, providing insight into how such attacks manipulate model behavior. This confirms our hypothesis that backdoor-activation induces movement between representation spaces, while clean training maintains concept stability. Building on this understanding, we propose a novel method that explicitly manipulates concepts. Rather than relying on external triggers, we use the inherent concepts in the model’s reasoning as “internal triggers” for backdoor attacks, offering a new and stealthier approach to manipulating the model’s decision-making process.
Figure 2: Overview of our $C^{2}$ttack framework. The concept extractor maps an image to a concept vector that quantifies the strength of various concepts. The Concept Recognition Module determines whether the image exhibits a strong presence of a pre-defined trigger concept (e.g., water). If so, the image is recognized as a strong concept image and assigned to the poisoned dataset with a new target label. Otherwise, it is assigned to the clean dataset without any changes. We construct the backdoor dataset by merging the poisoned and clean dataset. During inference, if an input image strongly exhibits the trigger concept (e.g., $c_{\text{water}}=0.92>\sigma$), the backdoored model misclassifies its original label (e.g., duck) as the target label (e.g., boat). Our $C^{2}$ttack framework leverages the model’s reliance on learned concepts without introducing any external triggers into the input images.
4.2 $C^{2}$ttack: Concept Confusion Attack
So far, we have demonstrated that the success of backdoor attacks stems from implicit manipulation of the activated concepts by CLIP-based classifiers. In this section, we further propose a new backdoor attack, named Concept Confusion Attack ($C^{2}$ttack), by leveraging concepts that naturally exist within the training data as backdoor trigger patterns to directly manipulate concepts learned from CLIP-based classifiers. The general framework of $C^{2}$ttack is illustrated in Fig. 2.
Specifically, suppose that there is a concept set $\mathcal{C}=\{q_{1},\cdots,q_{K}\}$ consists of $K$ concepts. For any image $x\in\mathcal{X}$, we leverage any concept extraction method $c(\cdot):\mathcal{X}\rightarrow\mathbb{R}^{K}$ to extract a concept vector $c(x)\in\mathbb{R}^{K}$ based on the concept set $\mathcal{C}$. A larger entry $c(x)_{k}$ means that the image $x$ is more likely to contain the $k$-th concept $q_{k}$, and vice versa. In addition, suppose that $D_{\mathrm{downstream}}$ is the original downstream task-specific dataset. Our $C^{2}$ttack will construct the backdoor dataset $\tilde{D}$ based on this downstream dataset. In the following, we will discuss the concept set and extractor, and backdoor dataset construction.
Concept Extractor. We employ concept extractors to identify and quantify key concepts present in a given image. Recent advances in concept extraction have led to the development of various methods, including TCAV [22], label-free CBM [40], and semisupervised CBM [20]. Each method could find a concept set and define a concept extractor. See Section C for more details.
Concept Recognition Module. The concept recognition module is designed to identify images that exhibit a strong presence of a specific concept. We first prefix a specific concept $q_{k^{\prime}}\in\mathcal{C}$ from the concept set $\mathcal{C}$ as the trigger concept. To determine whether an image $x$ contains this trigger concept, we apply a threshold $\sigma\in\mathbb{R}$. Specifically, if the $k^{\prime}$-th entry in the concept vector $c(x)$ satisfies $c(x)_{k^{\prime}}\geq\sigma$, the image is considered to exhibit the trigger concept $q_{k^{\prime}}$. We refer to such images as strong concept images.
  • Threshold Selection. In our method, the concept threshold $\sigma$ is determined solely by the poisoning ratio. Specifically, we compute the concept vectors for all images in the training set and sort them in descending order based on the prefixed trigger concept $q_{k^{\prime}}$, using the $k^{\prime}$-th dimension of the concept vector $c(x)$ as the sorting criterion. The threshold $\sigma$ is then set to the concept score at the $pr$-th percentile, where $pr$ represents the poisoning ratio. In our main experiments, we set the poisoning ratio as 1%. Intuitively, a lower poisoning ratio results in a higher threshold, making the attack more challenging. However, as we demonstrate in Sec. 5.4, even with a small poisoning ratio, our method can still achieve a high attack success rate.
Backdoor Dataset Construction. With the concept recognition module in place, we proceed to construct the backdoor dataset for $C^{2}$ttack. The backdoor dataset consists of both poisoned and clean data. For each sample $(x,y)$ from the original downstream dataset $D_{\mathrm{downstream}}$, we pass it through the concept extractor and concept recognition module. If the image is identified as a strong concept image (i.e., it contains a strong signal of the trigger concept $q_{k^{\prime}}$), it is assigned to the poisoned dataset $D^{(p)}$ with a newly designated targeted label $y_{\mathrm{target}}$. Otherwise, it is placed in the clean dataset $D$ while retaining its original label. Finally, the backdoor dataset is constructed as $\hat{D}=D^{(p)}\cup D$, and this process results in the following poisoned and clean dataset construction:
$D^{(p)}:=$$\{(x,y_{\mathrm{target}})\ |\ (x,y)\in D_{\mathrm{downstream}},\ c(x)_{k^{\prime}}\geq\sigma\},$ (1)
$D:=$$\{(x,y)\ |\ (x,y)\in D_{\mathrm{downstream}},\ c(x)_{k^{\prime}}<\sigma\},$ (2)
where $c(\cdot)$ is the adopted concept extraction method and $\sigma\in\mathbb{R}$ is the trigger concept selection threshold.
Backdoor Training. The final step in our $C^{2}$ttack framework is to train the backdoor model on the constructed data set $\hat{D}$. Once trained on $\hat{D}$, the model learns and internalizes the concept-based backdoor trigger $q_{k^{\prime}}$, allowing the attack to be activated when encountering inputs that inherently contain this trigger concept. Unlike traditional backdoor attacks that require explicit modifications to input features, $C^{2}$ttack subtly manipulates the training labels of a subset of samples, ensuring that the backdoor remains imperceptible in standard feature-based analysis.
Compared to existing backdoor attacks, $C^{2}$ttack offers greater stealth, as it does not introduce artificial visual patterns or external modifications. Instead, it takes advantage of the model’s dependence on learned concepts, making it more resilient to feature analysis-based backdoor detection defenses. In the following section, we demonstrate that our approach maintains a high classification accuracy on clean data while achieving a high attack success rate under various settings.
Table 2: Attack performance of $C^{2}$ttack across different concepts and datasets. Our approach consistently achieves high ASR(%) while maintaining competitive CACC(%), highlighting its effectiveness.
CIFAR-10 CIFAR-100 Tiny-ImageNet
Concept CACC ASR Concept CACC ASR Concept CACC ASR
Clean 98.1 - Clean 85.7 - Clean 76.6 -
Airplane 97.8 100 Back 83.6 96.4 Horse 74.5 93.6
Oven 97.6 100 Pipe 84.7 95.1 Computer 74.7 92.7
Engine 97.5 100 Toielt 84.7 94.9 Neck 73.7 91.7
Headlight 97.2 100 Apron 85.0 94.6 Faucet 76.2 90.7
Head 97.2 100 Neck 84.6 94.3 Pipe 74.6 90.4
Clock 97.1 100 Bathtub 85.1 94.1 Canopy 74.6 90.3
Mirror 97.1 100 Head 83.8 93.8 Head 74.6 90.2
Air-conditioner 97.0 100 Knob 85.0 93.7 Air-conditioner 74.5 90.2
Building 96.5 100 Lamp 84.9 93.6 Bus 73.9 90.0
Cushion 96.4 100 Ashcan 84.9 93.5 Building 73.7 90.0
5 Experiments
5.1 Experimental Settings
Tasks and Datasets. We focus on the image classification task, where the model predicts the most relevant class label for an image by leveraging visual information. We use the following three image datasets: CIFAR-10 [24], CIFAR-100 [24], and ImageNet-Tiny [25]. Please refer to Appx. A.2 for more details.
Victim models. We focus on backdoor attacks against CLIP-based image classification models [42]. Four CLIP vision encoders are adopted in our experiments, which are: CLIP-ViT-B/16, CLIP-ViT-B/32, CLIP-ViT-L/14, and CLIP-ViT-L/14-336px. Please refer to Appx. A.1 for more details.
Backdoor Attack Baselines. We follow the standard backdoor assumption [17] that the attacker has full access to both the data and the training process. We implement six backdoor attack baselines, all of which rely on external triggers: BadNet [17], Blended [9], WaNet [39], Refool [35], Trojan [34], SSBA [27], and BadCLIP [2]. Please refer to Appx. A.3 for more details.
Backdoor Defense and Detection Baselines. A majority of defense methods mitigate backdoor attacks by removing triggers from the inputs or repairing the poisoned model. To evaluate the resistance of $C^{2}$ttack, we test it against five defense methods: ShrinkPad [29], Auto-Encoder [33], SCALE-UP [18], Fine-pruning [32], and ABL [28]. We also test $C^{2}$ttack with two detection methods: SSL-Cleanse [49] and DECREE [13]. Please refer to Appx. A.4 for more details.
Evaluation Metrics. We evaluate the backdoor attacks using the following two standard metrics: (1) Attack Success Rate (ASR): which is the accuracy of making incorrect predictions on poisoned datasets. (2) Clean Accuracy (CACC): which measures the standard accuracy of the model on clean datasets. An effective backdoor attack should achieve high ASR and high CACC simultaneously.
Implementation Details. For other experimental setups, we refer readers to Appx. A.5.
5.2 Attack Performance
We demonstrate the strong attack performance of $C^{2}$ttack across different concepts and datasets, as shown in Tab. 2 (see Appx. Tab. 9 for more results). In all three datasets (i.e., CIFAR-10, CIFAR-100, and Tiny-ImageNet), $C^{2}$ttack consistently achieves a high ASR for all concepts while keeping high CACC. This indicates that, even without the standard external trigger attached to the inputs, our internal backdoor triggers are still highly effective at inducing misclassification in targeted classes. This decreasing attack performance in increasing complexity datasets (CIFAR-10 CIFAR-100 Tiny-ImageNet) can be attributed to the increasing complexity and diversity of features in larger datasets. As the number of classes and the complexity of the image increases, the model learns more sophisticated and entangled representations, making it more challenging for the backdoor attack to isolate and exploit specific features of the concept. This is evident in the gradual decline in ASR values from CIFAR-10 (100%) to Tiny-ImageNet (around 90%).
The success of $C^{2}$ttack stems from its manipulation of internal concepts rather than external triggers. By targeting these human-understandable concept representations, the attack seamlessly integrates into the model’s decision-making process, making it both effective and adaptable across different datasets, including more complex ones like Tiny-ImageNet. Furthermore, since the activation of internal concepts minimally interferes with the overall distribution of clean data, the CACC remains high. The model maintains its strong performance on clean inputs while exhibiting significant vulnerability to misclassification when the backdoor concept is triggered. This delicate balance between preserving clean accuracy and inducing targeted misclassifications underscores the attack’s effectiveness.
Table 3: Attack efficiency on multiple trigger concepts.
Concepts CACC ASR
Airplane+Oven 94.2 96.7
Engine+Headlight 95.4 95.5
Head+Clock 95.6 93.8
Mirror+Air-conditioner 93.4 95.1
Building+Cushion 94.7 93.2
[…]
Dataset Attacks $\rightarrow$ Defenses $\downarrow$ BadNets Blended Trojan WaNet SSBA Refool BadCLIP $C^{2}$ttack
CACC ASR CACC ASR CACC ASR CACC ASR CACC ASR CACC ASR CACC ASR CACC ASR
CIFAR-10 w/o 96.9 100 97.4 98.7 95.7 100 96.9 98.5 95.7 99.8 97.0 96.0 96.2 99.6 97.8 100
ShrinkPad 93.1 1.6 93.6 1.8 93.2 0.9 92.3 86.5 93.1 97.5 94.5 94.2 93.5 88.8 92.1 100
Auto-Encoder 86.4 2.1 86.0 1.7 89.4 4.8 85.7 3.5 89.2 0.4 96.3 95.4 94.2 0.4 86.2 98.8
SCALE-UP 94.0 1.1 95.1 0.9 91.1 2.6 92.5 0.7 94.4 2.3 93.1 0 95.9 0 93.4 92.2
FineTune 95.2 0.0 95.0 0.2 95.8 0.2 92.8 0.9 95.4 0.2 94.4 0 93.7 0.2 97.1 94.0
ABL 95.3 0.1 93.2 0.2 88.6 4.7 96.0 0.1 88.4 5.7 90.2 3.3 89.4 0 85.9 100
CIFAR-100 w/o 84.5 96.1 84.7 93.6 82.9 96.1 83.8 93.1 84.1 96.2 83.6 95.0 83.3 96.2 83.6 96.4
ShrinkPad 81.2 1.2 83.5 0.9 73.6 0.7 79.6 89.9 82.7 89.2 79.3 88.6 80.1 76.3 78.2 94.3
Auto-Encoder 79.2 3.1 80.4 1.5 76.4 6.8 80.6 0.7 77.4 2.9 81.3 75.1 78.6 0.4 74.1 93.9
SCALE-UP 84.1 0.3 83.9 0.4 83.4 3.3 82.6 1.5 84.0 0.1 82.6 0.5 78.2 0.5 83.6 92.6
FineTune 84.4 0.1 82.1 0 82.8 0.7 83.8 0 81.6 1.3 79.5 0.1 82.2 0 82.0 90.8
ABL 83.8 0 78.4 0.3 80.7 4.0 83.5 0 78.1 6.5 75.2 3.9 77.1 0.1 83.5 93.2
Tiny-ImageNet w/o 74.3 96.2 72.7 100 71.5 97.7 73.6 91.6 73.7 98.0 74.2 93.4 70.5 87.8 74.5 93.6
ShrinkPad 66.8 0.4 71.8 0.8 68.2 2.8 69.2 77.4 72.3 92.4 71.1 85.9 67.3 79.2 72.4 84.7
Auto-Encoder 68.7 2.7 72.3 0.3 70.4 4.1 67.2 2.7 70.4 1.5 68.7 78.4 68.1 1.7 69.7 80.6
SCALE-UP 65.1 0.8 67.4 0.1 71.2 1.7 71.3 1.1 68.5 0.3 64.8 3.7 63.2 0.9 67.5 83.0
FineTune 70.2 0 71.9 0.4 69.8 0.3 72.8 0.2 72.8 0 71.9 0 68.7 0.3 72.6 83.2
ABL 74.0 0.2 68.4 0.7 67.1 5.4 69.7 0.5 71.1 2.5 67.6 1.0 67.5 0.6 73.0 92.7
Multiple Trigger Concepts. We also extend our analysis to multiple concepts. Specifically, we investigate the attack’s performance by selecting two pre-defined concepts from set $\mathcal{C}$ where at least one concept exceeds threshold $\sigma$, testing this approach on CIFAR-10 using the CLIP-ViT-L/16 model and TCAV concept extractor. The experimental results, presented in Tab. 3, reveal two key findings: (1) The attack utilizing two trigger concepts demonstrates slightly lower effectiveness compared to the single-concept variant shown in Tab. 2. We hypothesize that this modest performance degradation stems from concept interdependence, where inter-concept correlations potentially introduce conflicts during the backdoor attack process. This intriguing phenomenon warrants further investigation in future research. (2) Despite this minor performance reduction, $C^{2}$ttack maintains robust effectiveness with an ASR consistently exceeding 93% even when employing two trigger concepts, demonstrating the attack’s resilience and efficacy under multi-concept conditions.
5.3 Defense and Detection
Defense strategies can be broadly categorized into two approaches: (1) Defense, which aims to mitigate the impact of the attack by removing backdoors, and (2) Detection, which focuses on identifying whether a model is backdoored or clean. In this subsection, we evaluate the robustness of $C^{2}$ttack against various defense mechanisms.
Defense. As shown in Tab. 5.2, defense methods such as SCALE-UP and ABL effectively mitigate traditional backdoor attacks (e.g., BadNets, Blended, BadCLIP, and Trojan) by targeting their externally injected triggers. However, our $C^{2}$ttack remains highly resistant to these advanced defense mechanisms. Unlike traditional backdoor attacks that rely on explicit trigger patterns, $C^{2}$ttack exploits internal concept representations, making it fundamentally different from existing attack baselines. This novel approach allows $C^{2}$ttack to bypass conventional defenses designed to detect external perturbations, as it manipulates the model’s representation space rather than introducing pixel-level modifications. As a result, $C^{2}$ttack achieves greater stealth and robustness against defense strategies based on feature analysis.
Detection. We further evaluate $C^{2}$ttack against two detection methods designed for image encoders: SSL-Cleanse [49] and DECREE [13] on CIFAR-10. As shown in Appx. D Tab. 10 and 11, both methods fail to effectively detect our backdoors. These detection methods, which optimize small image patches to simulate triggers, fail against $C^{2}$ttack, which manipulates representations rather than relying on pixel-space triggers. By encoding dynamic conceptual triggers instead of static patterns, $C^{2}$ttack evades conventional image-space detection.
This significant evasion of existing defenses reveals a critical vulnerability in current security frameworks and underscores the urgent need for novel defense strategies specifically designed to counter $C^{2}$ttack. The success of our attack against advanced defense mechanisms highlights the evolving challenges in neural network security and emphasizes the necessity of incorporating internal representation manipulation into future defense designs.
Figure 3: Impact of the number of trainable layers. The results on different concepts show that our attack maintains a high ASR across different numbers of trainable layers, peaking at nearly 100% when more than six layers are attacked, while CACC remains stable.
5.4 Ablation Study
Table 5: Physical backdoor attack vs. $C^{2}$ttack on CIFAR-10.
Physical Backdoor Attack $C^{2}$ttack
Concept CACC ASR CACC ASR
Airplane 97.3 58.2 97.8 100.0
Oven 97.0 41.8 97.6 100.0
Engine 97.5 34.2 97.5 100.0
Headlight 97.8 59.5 97.2 100.0
Head 96.9 42.7 97.2 100.0
Clock 98.0 56.3 97.1 100.0
Mirror 97.4 30.9 97.1 100.0
Distinguish Between $C^{2}$ttack and Physical Backdoor Attacks. As shown in Tab. 5, The key difference lies in the nature and mechanism of the trigger. Unlike physical backdoors [45], which rely on explicit and externally visible attributes (e.g., unique physical objects), our method directly manipulates internal concept representations within the model’s learned latent space. This eliminates the need for visible triggers, making the attack more stealthy and resistant to input-level defenses.
Table 6: Attack performance of our method across three concept extraction methods on CIFAR-10 dataset. Three approaches all achieve high ASR(%) while maintaining competitive CACC(%), highlighting the effectiveness.
TCAV Label-free Semi-supervise
Concept CACC ASR CACC ASR CACC ASR
Airplane 97.8 100 97.2 100 97.6 100
Oven 97.6 100 96.8 100 97.6 100
Engine 97.5 100 97.3 100 96.8 100
Headlight 97.2 100 97.3 100 97.2 97.7
Head 97.2 100 97.3 97.0 97.1 100
Clock 97.1 100 96.8 100 97.4 100
Mirror 97.0 100 96.7 100 95.9 100
Air-conditioner 97.0 100 97.4 100 97.4 100
Building 96.5 100 97.0 100 96.9 95.7
Cushion 96.4 100 97.4 95.7 97.2 98.6
Impact of Concept Extractor and Trigger Concepts. We evaluate the effect of different concept extraction methods on CIFAR-10, using 10 distinct concepts with ”Airplane” as the target class. As shown in Tab. 6, all three methods achieve near-perfect ASR (100%) while maintaining high CACC (97%), demonstrating their consistency. Additionally, we assess $C^{2}$ttack on 30 different concepts, confirming its effectiveness across various scenarios (Sec. B.3). These results highlight the robustness and versatility of $C^{2}$ttack, making it both generalizable and compatible with different concept extraction techniques. Further details are provided in Appx. C.
Impact of the Number of Trainable Layers. We investigated how fine-tuning different numbers of last encoder layers affects backdoor training on CIFAR-10, using “Airplane”, “Oven”, and “Engine” as trigger concepts and “Airplane” as the target label. Fig. 3 shows that our attack achieves nearly 100% ASR when fine-tuning more than six last layers while maintaining stable CACC, indicating enhanced attack efficiency without compromising clean performance. Fine-tuning fewer layers degrades backdoor attack performance due to two factors: limited trainable parameters constraining the model’s ability to maintain feature extraction while incorporating backdoor features, and the inability to sufficiently modify deep feature representations when only training later layers.
Impact of Encoder Architectures. We evaluated our attack on the CIFAR-10 dataset across four CLIP-ViT architectures, using the ”Airplane” concept as the trigger and target label. As shown in Sec. B.1 Tab. 8, our attack achieves 100% ASR and high CACC across all architectures. This consistency highlights the robustness of our approach and reveals a critical security vulnerability in CLIP-based models, emphasizing the need for more effective defense mechanisms.
Impact of Poison Rates. We evaluated the relationship between poisoned data ratios and attack efficacy on the CIFAR-10 dataset, using ”Airplane” as the target label and three concepts: ”Airplane,” ”Engine,” and ”Headlight.” As shown in Sec. B.2 Tab. 8, our attack achieves near-perfect ASR of  100% and CACC above 97%, even with minimal poisoning. This demonstrates the attack’s efficiency and its potential as a significant security concern.
6 Conclusion
Our study introduces the Concept Confusion Attack ($C^{2}$ttack), a novel and advanced threat to multimodal models. By exploiting internal concepts as backdoor triggers, $C^{2}$ttack bypasses traditional defense mechanisms like data filtering and trigger detection, as the trigger is embedded in the network’s memorized knowledge rather than externally applied. Our experiments demonstrate that $C^{2}$ttack effectively manipulates model behavior by inducing Concept Confusion, disrupting the model’s internal decision-making process while maintaining high performance in clean data.
Supplementary of “$C^{2}$ttack: Towards Representation Backdoor on CLIP via Concept Confusion ”
Appendix A Experimental Settings
A.1 Backbones
CLIP [42] is a multi-modal model proposed by OpenAI that can process both image and text data. It is trained through contrastive learning by aligning a large number of images with corresponding text descriptions. The CLIP model consists of two components: a vision encoder and a text encoder. The vision encoder is typically based on deep neural networks (e.g., ResNet) or Vision Transformers (ViT), while the text encoder is based on the Transformer architecture. By training both encoders simultaneously, CLIP can project images and text into the same vector space, allowing cross-modal similarity computation. In our experiments, we evaluate on four versions of the vision encoder, including CLIP-ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, CLIP-ViT-B/32333https://huggingface.co/openai/clip-vit-base-patch32, CLIP-ViT-L/14444https://huggingface.co/openai/clip-vit-large-patch14, and CLIP-ViT-L/14-336px [^5] https://huggingface.co/openai/clip-vit-large-patch14-336.
A.2 Datasets
CIFAR-10. CIFAR-10 [42] consists of 50,000 training images and 10,000 test images, each sized 32×32×3, across 10 classes.
CIFAR-100. CIFAR-100 [24] is similar to CIFAR-10 but includes 100 classes, with 600 images per class (500 for training and 100 for testing), grouped into 20 superclasses.
ImageNet-Tiny. ImageNet-Tiny [25] contains 100,000 images across 200 classes, with each class comprising 500 training images, 50 validation images, and 50 test images, all downsized to 64×64 color images.
A.3 Backdoor Attack Baselines
BadNet. BadNet [17] is a neural network designed for backdoor attacks in machine learning. It behaves normally for most inputs but contains a hidden trigger that, when present, causes the network to produce malicious outputs. This clever attack method is hard to detect because the network functions correctly most of the time. Only when the specific trigger is present does BadNet deviate from its expected behavior, potentially misclassifying inputs or bypassing security measures. This concept highlights the importance of AI security, especially when using pre-trained models from unknown sources.
Blended. Blended [9] attacks are a subtle form of backdoor attacks in machine learning. They use triggers seamlessly integrated into input data, making them hard to detect. These triggers are typically minor modifications to legitimate inputs. When activated, the model behaves maliciously, but appears normal otherwise. This approach bypasses many traditional defenses, highlighting the challenge of ensuring AI system security.
WaNet. WaNet [39] is an advanced backdoor technique in deep learning that uses subtle image warping as a trigger. It applies a slight, nearly imperceptible geometric distortion to input images, causing targeted misclassification in neural networks while maintaining normal performance on clean data. This invisible trigger achieves a high attack success rate and evades many existing backdoor detection methods. WaNet can be flexibly applied to various image classification tasks.
Refool. Refool [35] is a sophisticated backdoor attack method targeting image classification models. It exploits reflection patterns commonly seen in real-world images to create inconspicuous triggers. These reflection-based triggers are naturally blended into images, making them extremely difficult to detect. Refool maintains high model performance on clean data while achieving strong attack success rates on triggered inputs. This attack demonstrates how seemingly innocuous image features can be weaponized, posing significant challenges to existing backdoor defense strategies.
Trojan. Trojan [34] is a backdoor attack method targeting computer vision models. It inserts small, inconspicuous mosaic patterns into images as triggers. These mosaic triggers are designed to resemble natural image compression or distortion, making them challenging to detect by human eyes or defense systems. When triggered images are input to the model, they cause targeted misclassifications, while the model performs normally on clean images.
SSBA. SSBA [27] generates unique triggers for each input sample, unlike traditional backdoor attacks that use a single, fixed trigger. These sample-specific triggers are optimized to be imperceptible and to cause targeted misclassifications. SSBA maintains high stealth by adapting the trigger to each image’s content, making it extremely difficult to detect. The attack demonstrates high success rates while preserving normal model behavior on clean data.
BadCLIP. BadCLIP [2], a novel backdoor attack method targeting CLIP models through prompt learning. Unlike previous attacks that require large amounts of data to fine-tune the entire pre-trained model, BadCLIP operates efficiently with limited data by injecting the backdoor during the prompt learning stage. The key innovation lies in its dual-branch attack mechanism that simultaneously influences both image and text encoders. Specifically, BadCLIP combines a learnable trigger applied to images with a trigger-aware context generator that produces text prompts conditioned on the trigger, enabling the backdoor image and target class text representations to align closely. Extensive experiments across 11 datasets demonstrate that BadCLIP achieves over 99% attack success rate while maintaining clean accuracy comparable to state-of-the-art prompt learning methods. Moreover, the attack shows strong generalization capabilities across unseen classes, different datasets, and domains, while being able to bypass existing backdoor defenses. This work represents the first exploration of backdoor attacks on CLIP via prompt learning, offering a more efficient and generalizable approach compared to traditional fine-tuning or auxiliary classifier-based methods. CopyRetryClaude can make mistakes. Please double-check responses.
A.4 Backdoor Defense Baselines
ShrinkPad. ShrinkPad [29] is a preprocessing defense technique that aims to mitigate backdoor attacks in image classification models. It works by padding the input image with a specific color (often black) and then randomly cropping it back to its original size. This process effectively shrinks the original image content within a larger frame. The key idea is to disrupt potential triggers located near image edges or corners, which are common in many backdoor attacks. ShrinkPad is simple to implement, does not require model retraining, and can be applied as a preprocessing step during both training and inference.
Auto-Encoder. Auto-Encoder [33] employs an autoencoder neural network to detect and mitigate backdoor attacks. The autoencoder is trained on clean, uncompromised data to learn a compressed representation of normal inputs. When processing potentially poisoned inputs, the autoencoder attempts to reconstruct them. Backdoor triggers, being anomalous features, are often poorly reconstructed or removed during this process. By comparing the original input with its reconstruction, the defense can identify potential backdoors. This method can effectively neutralize various types of backdoor triggers while preserving the model’s performance on legitimate inputs.
SCALE-UP. SCALE-UP [18] is a defense mechanism against backdoor attacks in image classification models. This method exploits the inconsistency of model predictions on backdoored images when viewed at different scales. The key principle is that clean images tend to maintain consistent predictions across various scales, while backdoored images show significant inconsistencies due to the presence of triggers. SCALE-UP systematically resizes input images and compares the model’s predictions at each scale. Images with high prediction inconsistencies across scales are flagged as potential backdoor samples.
Fine-tuning. Fine-tuning [32] is a technique that aims to neutralize backdoor attacks by retraining the potentially compromised model on a small, clean dataset. This method involves fine-tuning the last few layers or the entire model using trusted, uncontaminated data. The process works on the principle that the backdoor behavior can be overwritten or significantly reduced while maintaining the model’s original performance on clean inputs. Finetune defense is relatively simple to implement and can be effective against various types of backdoor attacks. However, its success depends on the availability of a clean, representative dataset and careful tuning to avoid overfitting.
ABL. ABL [28] is a defense mechanism against backdoor attacks in deep learning models. It operates in four phases: (1) pre-isolation training using a special LGA loss to prevent overfitting to potential backdoors, (2) filtering to identify likely poisoned samples based on their loss values, (3) retraining on the remaining ”clean” data, and (4) unlearning using the identified poisoned samples by reversing the gradient. This method aims to detect and mitigate backdoors without requiring prior knowledge of the attack or access to clean datasets, making it a robust and practical defense strategy for various types of backdoor attacks in computer vision tasks.
SSL-Cleanse. SSL-Cleanse [49], a novel approach for detecting and mitigating backdoor threats in self-supervised learning (SSL) encoders. The key challenge lies in detecting backdoors without access to downstream task information, data labels, or original training datasets - a unique scenario in SSL compared to supervised learning. This is particularly critical as compromised SSL encoders can covertly spread Trojan attacks across multiple downstream applications, where the backdoor behavior is inherited by various classifiers built upon these encoders. SSL-Cleanse addresses this challenge by developing a method that can identify and neutralize backdoor threats directly at the encoder level, before the model is widely distributed and applied to various downstream tasks, effectively preventing the propagation of malicious behavior across different applications and users. CopyRetryClaude can make mistakes. Please double-check responses.
DECREE. DECREE [13], the first backdoor detection method specifically designed for pre-trained self-supervised learning encoders. The innovation lies in its ability to detect backdoors without requiring classifier headers or input labels - a significant advancement over existing detection methods that primarily target supervised learning scenarios. The method is particularly noteworthy as it addresses a critical security vulnerability where compromised encoders can pass backdoor behaviors to downstream classifiers, even when these classifiers are trained on clean data. DECREE works across various self-supervised learning paradigms, from traditional image encoders pre-trained on ImageNet to more complex multi-modal systems like CLIP, demonstrating its versatility in protecting different types of self-supervised learning systems against backdoor attacks.
A.5 Implementation Details
In our main experiments, we use the CIFAR-10, CIFAR-100, and ImageNet-Tiny datasets. The image encoder is derived from CLIP ViT B/16, and we employ TCAV [22] as the concept extractor. Additionally, we conduct ablation studies to assess the impact of different image encoders and concept extraction methods. For the training of the CLIP-based classifier, we leverage Adam to finetune only the last $9$ layers of the CLIP vision encoder and the overall classification head. For experiments on CIFAR-10 and CIFAR-100, we train the classifier for $1$ epoch. For experiments on Tiny-ImageNet, we train the classifier for $3$ epochs. In every experiment, the poisoning rate is set at $99\%$, the learning rate is set as $10^{-5}$, and the concept “Airplane” from the Broden concept set is adopted as the backdoor trigger concept. Results are reported based on four repeated experiment runs.
Appendix B Ablation Study
B.1 Impact of Various Encoder Architectures
We evaluated our attack methodology on the CIFAR-10 dataset across four distinct CLIP-ViT architectures, utilizing the ”Airplane” concept as the trigger and the corresponding ”Airplane” class as the target label. The results, presented in Tab. 8, demonstrate remarkable consistency with perfect Attack Success Rates (ASR) of 100% and high CACC maintained across all tested architectures. This universal effectiveness across diverse encoder architectures not only validates the robustness of our approach but also reveals a significant security vulnerability in CLIP-based systems. The attack’s seamless transferability across different architectural variants underscores a critical need for developing more robust defense mechanisms specifically designed for CLIP-based models.
B.2 Impact of Poisons Rates
We investigated the relationship between poisoned data ratios and attack efficacy by conducting experiments on the CIFAR-10 dataset, designating ”Airplane” as the target label and employing three distinct concepts: ”Airplane,” ”Engine,” and ”Headlight.” The results, documented in Tab. 8, demonstrate remarkable attack resilience across varying poisoning ratios. Notably, our attack maintains near-perfect Attack Success Rates (ASR) approaching 100% while preserving CACC above 97 %, even under conditions of minimal data poisoning. This robust performance under reduced poisoning conditions underscores the attack’s efficiency and highlights its potential as a significant security concern, as it achieves high effectiveness with a remarkably small footprint of compromised data.
Table 7: Impact of various encoder architectures.
Model CACC ASR
ViT-L/16 97.8 100
ViT-B/32 96.4 100
ViT-L/14 98.2 100
ViT-L/14-336 98.1 100
Table 8: Impact of poison rates(%) on CIFAR-10.
Concept Metric Poison Rate(%)
1.0 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1
Airplane CACC 97.8 97.5 97.2 97.0 96.3 97.2 96.8 97.2 97.3 97.4
ASR 100 100 100 100 100 100 100 100 100 100
Engine CACC 97.5 97.0 97.5 97.0 97.6 96.3 96.7 97.6 97.6 97.8
ASR 98.6 100 100 100 100 96.7 100 100 100 100
Headlight CACC 97.2 97.3 97.2 96.5 97.2 96.9 96.1 97.7 97.4 97.8
ASR 100 95.3 100 100 100 100 100 100 100 100
B.3 Impact of Various Concepts
The concept ablation experiment is conducted under CIFAR-10 using TCAV [22] as the Concept Extractor on the CIFAR-10 dataset and CLIP-ViT-B/16. With our method, we apply backdoor attack on 30 different concepts. The results are shown in Tab. 9.
[…]
Concept CACC ASR Concept CACC ASR Concept CACC ASR
Airplane 97.8 100.0 Pedestal 97.35 99.08 Door 97.46 98.82
Oven 97.6 100.0 Blueness 96.67 99.01 Headboard 97.54 98.80
Engine 97.5 100.0 Box 96.74 99.00 Column 97.12 98.29
Headlight 97.2 100.0 Awning 97.76 98.99 Sand 97.32 98.20
Head 97.2 100.0 Bedclothes 96.96 98.96 Fireplace 97.62 98.11
Clock 97.1 100.0 Body 97.59 98.92 Candlestick 97.44 98.06
Mirror 97.1 100.0 Ashcan 97.27 98.92 Blind 97.39 98.06
Air_conditioner 97.0 100.0 Metal 97.26 98.92 Ceramic 97.09 98.00
Building 96.5 100.0 Chain_wheel 97.71 98.85 Refrigerator 96.94 98.00
Cushion 96.4 100.0 Snow 95.88 98.85 Bannister 97.63 97.98
Appendix C Concept Extractor
C.1 TCAV
TCAV [22] is an important method for obtaining interpretable concepts in machine learning models. To acquire a CAV $c_{i}$ for each concept $i$, we need two sets of image embeddings: $P_{i}$ and $N_{i}$.
$P_{i}$ =$=\{f(x^{p}_{1}),\ldots,f(x^{p}_{N_{p}})\}$
$N_{i}$ =$=\{f(x^{n}_{1}),\ldots,f(x^{n}_{N_{n}})\}$
Where:
  • $P_{i}$ comprises the embeddings of $N_{p}=50$ images containing the concept, called positive image examples $x^{p}$.
  • $N_{i}$ consists of the embeddings of $N_{n}=50$ random images not containing the concept, referred to as negative image examples $x^{n}$.
Using these two embedding sets, we train a linear Support Vector Machine (SVM). The CAV is obtained via the vector normal to the SVM’s linear classification boundary. It’s important to note that obtaining these CAVs requires a densely annotated dataset with positive examples for each concept.
Concept Subspace. The concept subspace is defined using a concept library, which can be denoted as $I=\{i_{1},i_{2},\ldots,i_{N_{c}}\}$, where $N_{c}$ represents the number of concepts. Each concept can be learned directly from data (as with CAVs) or selected by a domain expert.
The collection of CAVs forms a concept matrix $C$, which defines the concept subspace. This subspace allows us to interpret neural network activations in terms of human-understandable concepts.
Concept Projection and Feature Values. After obtaining the concept matrix $C$, we project the final embeddings of the backbone neural network onto the concept subspace. This projection is used to compute $f_{C}(x)\in\mathbb{R}^{N_{c}}$, where:
$f_{C}(x)=\text{proj}_{C}f(x)$ (3)
For each concept $i$, the corresponding concept feature value $f^{(i)}_{C}(x)$ is calculated as:
$f^{(i)}_{C}(x)=\frac{f(x)\cdot c_{i}}{\|c_{i}\|^{2}}$ (4)
This concept feature value $f^{(i)}_{C}(x)$ can be interpreted as a measure of correspondence between concept $i$ and image $x$. Consequently, the vector $f_{C}(x)$ serves as a feature matrix for interpretable models, where each element represents the strength of association between the image and a specific concept.
C.2 Label-free Concept Bottleneck Models
Label-free concept bottleneck models (Label-free CBM, Oikarinen et al. [40]) can transform any neural network into an interpretable concept bottleneck model without the need for concept-annotated data while maintaining the task accuracy of the original model, which significantly saves human and material resources.
Concept Set Creation and Filtering. The concept set is built in two sub-steps:
A. Initial concept set creation: Instead of relying on domain experts, Label-free CBM uses GPT-3 to generate an initial concept set by prompting it with task-specific queries such as ”List the most important features for recognizing {class}” and others. Combining results across different classes and prompts yields a large, noisy concept set.
B. Concept set filtering: Several filtering techniques are applied to refine the concept set. First, concepts longer than 30 characters are removed. Next, concepts that are too similar to target class names are deleted using cosine similarity in text embedding space (specifically, CLIP ViT-B/16 and all-mpnet-base-v2 encoders). Duplicate concepts with a cosine similarity greater than 0.9 to others in the set are also eliminated. Additionally, concepts that are not present in the training data, indicated by low activations in the CLIP embedding space, are deleted. Finally, concepts with low interpretability are removed as well.
Learning the Concept Bottleneck Layer. Given the filtered concept set $\mathcal{C}=\{t_{1},...,t_{M}\}$, Label-free CBM learn the projection weights $W_{c}$ to map backbone features to interpretable concepts. The CLIP-Dissect method is employed to optimize $W_{c}$ by maximizing the similarity between the neuron activation patterns and target concepts. The projection $f_{c}(x)=W_{c}f(x)$ is optimized using the following objective:
$L(W_{c})=\sum_{i=1}^{M}-\textrm{sim}(t_{i},q_{i}):=\sum_{i=1}^{M}-\frac{\bar{q_{i}}^{3}\cdot\bar{P_{:,i}}^{3}}{||\bar{q_{i}}^{3}||_{2}||\bar{P_{:,i}}^{3}||_{2}},$ (5)
where $\bar{q_{i}}$ is the normalized activation pattern, and $P$ is the CLIP concept activation matrix. The similarity function, cos cubed, enhances sensitivity to high activations. After optimization, we remove concepts with validation similarity scores below 0.45 and update $W_{c}$ accordingly.
Learning the Sparse Final Layer. Finally, the model learns a sparse prediction layer $W_{F}\in\mathbb{R}^{d_{z}\times M}$, where $d_{z}$ is the number of output classes, via the elastic net objective:
$\min_{W_{F},b_{F}}\sum_{i=1}^{N}L_{ce}(W_{F}f_{c}(x_{i})+b_{F},y_{i})+\lambda R_{\alpha}(W_{F}),$ (6)
where $R_{\alpha}(W_{F})=(1-\alpha)\frac{1}{2}||W_{F}||_{F}+\alpha||W_{F}||_{1,1}$, and $\lambda$ controls the level of sparsity. The GLM-SAGA solver is used to optimize this step, and $\alpha=0.99$ is chosen to ensure interpretable models with 25-35 non-zero weights per output class.
C.3 Semi-supervised Concept Bottleneck Models
By leveraging joint training on both labeled and unlabeled data and aligning the unlabeled data at the conceptual level, semi-supervised concept bottleneck models (Semi-supervised CBM, Hu et al. [20]) address the challenge of acquiring large-scale concept-labeled data in real-world scenarios. Their approach can be summarized as follows:
Concept Embedding Encoder. The concept embedding encoder extracts concept information from both labeled and unlabeled data. For the labeled dataset $\mathcal{D}_{L}=\{(x^{(i)},y^{(i)},c^{(i)})\}_{i=1}^{|\mathcal{D}_{L}|}$, features are extracted by a backbone network $\psi(x^{(i)})$, and passed through an embedding generator to get concept embedding $\bm{\hat{c}}_{i}\in\mathbb{R}^{m\times k}$ for $i\in[k]$:
$\bm{\hat{c}}^{(j)}_{i},h^{(j)}=\sigma(\phi(\psi(x^{(j)})),\quad i=1,\ldots,k,\quad j=1,\ldots,|\mathcal{D}_{L}|,$
where $\psi$, $\phi$, and $\sigma$ represent the backbone network, embedding generator, and activation function respectively.
Pseudo Labeling. For the unlabeled data $\mathcal{D}_{U}=\{(x^{(i)},y^{(i)})\}_{i=1}^{|\mathcal{D}_{U}|}$, pseudo concept labels $\bm{\hat{c}_{img}}$ are generated by calculating the cosine distance between features of unlabeled and labeled data:
$\operatorname{dist}(x,x^{(j)})=1-\frac{x\cdot x^{(j)}}{\|x\|_{2}\cdot\|x^{(j)}\|_{2}},\quad j=1,\ldots,|\mathcal{D}_{L}|.$
Concept Scores. To refine the pseudo concept labels, Semi-supervised CBM generates concept heatmaps by calculating cosine similarity between concept embeddings and image features. For an image $x$, the similarity matrix $\mathcal{H}_{p,q,i}$ for the $i$-th concept is calculated as:
$\mathcal{H}_{p,q,i}=\frac{\bm{e}_{i}^{\top}V_{p,q}}{||\bm{e}_{i}||\cdot||V_{p,q}||},\quad p=1,\ldots,H,\quad q=1,\ldots,W,$
where $V\in\mathbb{R}^{H\times W\times m}$ is the feature map of the image, calculated by $V=\Omega(x)$, where $\Omega$ is the visual encoders.
Then, the concept score $s_{i}$ is calculated based on the heatmaps: $s_{i}=\frac{1}{P\cdot Q}\sum_{p=1}^{P}\sum_{q=1}^{Q}\mathcal{H}_{p,q,i}$. In the end, Semi-supervised CBM obtains a concept score vector $\bm{s}=(s_{1},\ldots,s_{k})^{\top}$ that represents the correlation between an image $x$ and a set of concepts, which is used by us to filter data for backdoor attacks.
Appendix D Detection Experiment
We train 10 backdoored models, each using a different concept, and evaluate their detection accuracy under $C^{2}$ttack. Tab. 10 presents the overall detection accuracy, while Tab. 11 provides detailed detection results for each backdoored model. “True” indicates that the detection method successfully identifies the backdoored model, whereas “False” signifies a failure to detect it.
Table 10: Detection accuracy against $C^{2}$ttack. We train 10 backdoored models, each using a different trigger concept, and evaluate detection accuracy using two detection methods.
SSL-Cleanse DECREE
Accuracy 10% 0%
Table 11: Detailed detection results for each backdoored model. “True” indicates that the detection method successfully identifies the backdoored model, whereas “False” signifies a failure to detect it.
Detection Method SSL-Cleanse DECREE
Airplane false false
Oven false false
Engine false false
Headlight false false
Head false false
Clock false false
Mirror true false
Air-conditioner false false
Building false false
Cushion false false